* Question 4

On December 31, 2020, Blue Company acquired a computer from Plato Corporation by issuing a $609,000 zero-interest-bearing note, payable in full on December 31, 2024. Blue Company’s credit rating permits it to borrow funds from its several lines of credit at 12%. The computer is expected to have a 5-year life and a $63,000 salvage value.

Prepare the journal entry for the purchase on December 31, 2020. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answers to 0 decimal places e.g. 58,971. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

December 31, 2020

Prepare any necessary adjusting entries relative to depreciation (use straight-line) and amortization (use effective-interest method) on December 31, 2021. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

December 31, 2021

(To record the depreciation.)

December 31, 2021

(To amortize the discount.)

Schedule of Note Discount Amortization


Date

Debit, Interest Expense Credit,
Discount on Notes Payable

Carrying Amount
of Note

12/31/20 $

$

12/31/21

12/31/22

12/31/23

12/31/24

Prepare any necessary adjusting entries relative to depreciation and amortization on December 31, 2022. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

December 31, 2022

(To record the depreciation.)

December 31, 2022

(To amortize the discount.)
Answer #1

Solution 1:

Present value of note = $609,000 * PV factor at 12% for 4th period

= $609,000 * 0.63352 = $387,032

Date Account Titles and Explanation Debit Credit
31-Dec-20 //img.homeworklib.com/questions/39cbfde0-6fa6-11ea-aa99-05cc7aee8cad.gif?x-oss-process=image/resize,w_560
Computer Dr
$387,032.00
Discount on notes payable Dr $221,968.00
      To Notes Payable $609,000.00
(To record purchase of computer)

Solution 2:

Date Account Titles and Explanation Debit Credit
31-Dec-21 //img.homeworklib.com/questions/39cbfde0-6fa6-11ea-aa99-05cc7aee8cad.gif?x-oss-process=image/resize,w_560
Depreciation expense Dr [($387,032 - $63,000)/5]
$64,806.00
      To Accumulated depreciation $64,806.00
(To record depreciation)
31-Dec-21 Interest expense Dr $46,444.00
      To Discount on notes payable $46,444.00
(To amortize the discount)


Solution 3:

Schedule of Note Discount Amortization
Date Debit, Interest Expense Credit, Discount on Notes Payable Carrying Amount of Note
31-Dec-20 $387,032
31-Dec-21 $46,444 $433,476
31-Dec-22 //img.homeworklib.com/questions/39cbfde0-6fa6-11ea-aa99-05cc7aee8cad.gif?x-oss-process=image/resize,w_560
$52,017
//img.homeworklib.com/questions/39cbfde0-6fa6-11ea-aa99-05cc7aee8cad.gif?x-oss-process=image/resize,w_560
$485,493
31-Dec-23 $58,259 $543,752
31-Dec-24 $65,248 $609,000

Solution 4:

Date Account Titles and Explanation Debit Credit
31-Dec-22 //img.homeworklib.com/questions/39cbfde0-6fa6-11ea-aa99-05cc7aee8cad.gif?x-oss-process=image/resize,w_560
Depreciation expense Dr [($387,032 - $63,000)/5]
$64,806.00
      To Accumulated depreciation $64,806.00
(To record depreciation)
31-Dec-22 Interest expense Dr $52,017.00
      To Discount on notes payable $52,017.00
(To amortize the discount)